# Prompt Tuning Strikes Back: Customizing Foundation Models with Low-Rank Prompt Adaptation

### v1

@@ 1 Introduction
In recent years, numerous PEFT approaches have been proposed [11, 12, 16, 17, 18, 30, 36], each varying in how and what they modify within FMs [10, 9]. These variations can be categorized by the position in the FMs where modifications are applied, the functions used for modification, and the methods of integrating the modifications. While effective, these methods require maintaining multiple adapter-like modules for each task on the FM server and the need to select and assemble a subset of these modules every time a batch is processed for inference [35] (see Fig. 1).
@@
  • We verify its effectiveness by evaluating it against full fine-tuning and state-of-the-art PEFT methods on nine tasks using seven different transformer backbones.
@@ 3.1 Preliminaries
Transformers. A transformer model consists of multiple stacked layers, where each layer has multiple heads, each performing self-attention over the input [33]. Let us consider a single head, $H$ parameterised by the query, key, and value weights as $W^{Q},W^{K},W^{V}\in\mathcal{R}^{d_{H}\times d}$, respectively. For a given sequence of $n$ input vectors $\textbf{X}=\{\textbf{x}^{1},\ldots,\textbf{x}^{n}\}$ with each $\textbf{x}\in\mathcal{R}^{d}$ and a query vector $\textbf{x}^{i}$, the output of the head is
$\textbf{o}^{i}=\textrm{Attention}(W^{Q}\textbf{x}^{i},W^{K}\textbf{X},W^{V}\textbf{X})=\text{softmax}(W^{Q}\textbf{x}^{i}(W^{K}\textbf{X})^{\top})W^{V}\textbf{X}$ (1)
where, we have ignored the constant $\sqrt{d_{H}}$ for notational convenience. For the first layer, the input X is the embedding matrix $\textbf{X}_{e}\in\mathcal{R}^{d\times n}$.
@@
Prompt Tuning (PT). The objective of prompt tuning is to learn a "soft prompt" $\textbf{Z}\in\mathcal{R}^{d\times m}$ that is parameterised by a set of $m$ learnable vectors $\{\textbf{z}^{1},\ldots\textbf{z}^{m}\}$, where $m$ denotes its length [15]. They can be concatenated to the input word embeddings and trained via back-propagation: $\max_{\textbf{Z}}\mathcal{L}=\mathbb{E}[\log p_{W}(\textbf{y}|\mathrm{concat}(\textbf{Z},\textbf{X}_{e}))]$. The soft prompt can interpreted as an embedding of virtual tokens that enable the model to perform a downstream task without updating its parameters [28, 8].
@@
where, $A_{ik}$ is the attention transformer gives to the prefix vector $\textbf{z}_{k}$ for a given query vector $\textbf{x}_{i}$. We can observe that soft prompt linearly interpolates the head’s position-wise output. Thus, the bias term can be considered in an offset subspace spanned by vectors $\{W^{V}\textbf{z}^{k}\}$ with dimension $m$ (or $\leq m$) [10].
3.2 Our model
@@
Both $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$ have their own dedicated parameters. Similar to prompt tuning, $\textbf{Z}_{S}$ consists of $m$ learnable vectors. Conversely, $\textbf{Z}_{I}$ is obtained from input using the encoding function $f:\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times m}$. However, encoding a matrix of size $d\times m$ can be expensive. For example, in [38], an MLP layer with hidden dimension $h$ requires $\mathcal{O}(hdm)$ parameters. To improve parameter efficiency, we assume a low-rank decomposition of $\textbf{Z}_{I}$ and encode two matrices, $\textbf{u}\in\mathbb{R}^{d\times r}$ and $\textbf{v}\in\mathbb{R}^{m\times r}$, using separate MLP layers. This design makes $f$ computationally cheaper, with $\mathcal{O}(hdm(\frac{r}{d}+\frac{r}{m}))$ parameters, reducing the factor to $(\frac{r}{d}+\frac{r}{m})<1$ by choosing $r\ll\min\{d,m\}$. The overall composition of Z can be represented as:
@@
where _ marks the three trainable components. The MLP head consists of a down- and up-projection layer with a non-linear activation in between. It sits atop a smaller language model, referred to as the Encoder model, that gives the input embedding $\textbf{X}^{\prime}$. A visual illustration of the framework can be found in Fig. 2.
@@
Coupled learning of ${\textbf{Z}_{S}}$ and ${\textbf{Z}_{I}}$. Let’s examine the partial derivatives of $\mathcal{L}$ with respect to $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$. Using the chain-rule on the formulation in Eq. 4 with $g(.)=\textrm{sigmoid}(.)$, $\nabla_{\textbf{Z}_{S}}\mathcal{L}=\nabla_{\textbf{Z}}\mathcal{L}\circ\sigma(\textbf{Z}_{I})$, and $\nabla_{\textbf{Z}_{I}}\mathcal{L}=(\nabla_{\textbf{Z}}\mathcal{L}\circ\textbf{Z}_{S})\cdot\sigma(\textbf{Z}_{I})(1-\sigma(\textbf{Z}_{I}))$.
[…]
@@ 4.1 Experimental Setup
Tasks. We evaluate LOPA on (i) six natural language understanding tasks from the GLUE benchmark [34]—namely, SST-2 [31], MNLI [37], MRPC [3], QNLI [29], QQP, and RTE [5]; (ii) a code-generation task that requires the model to complete method bodies from MBPP benchmark [1], and (iii) two code-understanding tasks—namely, CruxEval-I (input prediction) and CruxEval-O (output prediction) from CruxEval benchmark [6]. These tasks assess the model’s ability to reason about the execution of simple Python programs by asking it to predict the input given the output of a function and vice-versa.
Baselines. We evaluate LOPA against five baselines that represent different customisation approaches. (1) FFT and (2) LoRA are selected as representatives of methods that customise models on the server side, with FFT representing full fine-tuning and LoRA [12] showcasing a parameter-efficient technique. (3) standard prompt tuning [15] and (4) S-IDPG[38] are chosen because they customise models from the user side, focusing on soft-prompt insertion. For IDPG, we use DNN layers in MLP-head to encode the soft prompt for the input embedding layer. We chose DNN over PHM layers [39], because we found them to exhibit better performance across our task set. Additionally, the proposed LOPA also uses DNN layers, facilitating a fair comparison with S-IDPG-DNN  [^2] For a further reduction in trainable parameters, LOPA can be used in conjunction with PHM layers.. Lastly, we include results for the model without any fine-tuning to report its (5) zero-shot performance.
Backbone Architectures. For NLU tasks, we implement all the tuning methods on 355M RoBERTa [22]. For Code Generation and Understanding, we extensively test the baselines on a range of FM backbones: starting from smaller FMs - 350M CodeGen-mono [25] and 1.3B Deepseek-Coder [7]; mid-sized FMS - 2.7B phi-2 [13], 3.8B phi-3 [24]; and larger FMs - 7B Deepseek-Coder [7] and 8B Llama 3 [23].
@@
Regarding evaluation, binary or multi-class classification accuracies and F1 scores, are employed for NLU tasks as provided in the GLUE benchmark. For coding tasks, we report the $\textit{pass}@1$ scores computed using the best-performing temperatures: $0.6$ for MBPP and $0.2$ for CruxEval.
@@ 4.2 Baseline Comparison
Performance on Natural Language Understanding. In Table 1, we can observe that the LOPA consistently outperforms the traditional prompt-tuning method by an average margin of $28.62$ points. This result demonstrates that conditioning the soft prompt on instances enables it to influence the model’s output more significantly. Furthermore, LOPA shows an average improvement of $5.36$ points over IDPG, highlighting that the proposed factorisation captures complex relationships between task-specific ($\textbf{Z}_{S}$) and instance-specific ($\textbf{Z}_{I}$) components. This improvement is particularly notable in limited-data settings, with a $12.5$-point increase in MRPC and a $6.13$-point increase in RTE.
@@
Performance on Code Understanding. In Table 2, LOPA consistently improves the $\textit{pass}@1$ score of the baseline with no tuning across all FM backbones. It outperforms prompt tuning (PT) on CruxEval tasks, with modest improvements of approximately $2$ to $4$ points on smaller FMs like CodeGen and DeepSeekCoder-1.3B and larger improvements ranging from $8$ to $11$ points on larger FMs like LLama-3 and Phi-3 in CruxEval-O. Furthermore, IDPG performs worse than PT for all models except Phi-3 in CruxEval-O. These results suggest that merely encoding an instance-sensitive soft prompt does not guarantee improvements and can even degrade performance (e.g., IDPG on CodeGen and DeepSeekCoder-1.3B in CruxEval tasks). The poor generalization of the learned soft prompt, possibly due to over-parameterization and resulting overfitting, might explain this behaviour. In contrast, LOPA, being more parameter-efficient, explicitly incorporates task and instance information in its design of the soft prompt, leading to better performance.
LOPA also performs on par with LoRA, often within a range of $1$ to $4$ points of $\textit{pass}@1$, while roughly using two-third the parameters. Notably, LOPA outperforms LoRA across all models in CruxEval-I, except for LLama-3, with improvements approximately ranging from $2$ points in DeepSeekCoder-7B to $4$ points in Phi-3. This result might be attributed to the fact that many of the FMs considered here are good knowledge approximators, well-trained on diverse datasets, and demonstrate strong zero-shot generalization. Directly updating their weights might lead to catastrophic forgetting, where the models lose previously acquired knowledge [27]. In such cases, soft prompting, as employed by LOPA, can effectively elicit the necessary skills to solve new tasks without compromising existing knowledge [26].
Performance on Code Completion. On the code completion task of MBPP, both IDPG and LOPA improve the performance of the baseline model with nearly equal gains except for LLama-3. However, LOPA achieves this with significantly fewer tunable parameters—approximately half the number used by CodeGen and only one-fifth of those used by IDPG in DeepseekCoder-7b. This demonstrates that LOPA scales well with the size of the foundation model, maintaining both performance and parameter efficiency. This efficiency is attributed to the low-rank approximation of the instance-specific matrix employed by LOPA. For LLama-3, all tuning approaches led to a drop in performance, possibly due to over-fitting, suggesting that LLama-3 might already be trained on MBPP.
@@ 4.3 Ablation Study
Performance as a function of soft-prompt length. In Figure 3, we study how the length of the soft prompt impacts the performance of LOPA compared to other prompting methods. Increasing the length of the soft prompt corresponds to adding more vectors to the set that represents the soft prompt, thus expanding the offset subspace (See Eq. 2). Whether the added vectors are mutually independent and provide additional useful information depends on the tuning approach to learning them and the offset subspace of the FM.
@@
Performance of LOPA as a function of rank. In Figure 4, we examine how the rank of the proposed low-rank decomposition of the instance-specific component affects the performance. For NLU, we consider three tasks: MRPC, RTE, and SST-2, and observe the performance of RoBERTa as the rank increases from 1 to 4 . Performance consistently improves with increasing rank, showing gains of $1$ to $2\%$ for SST-2 and up to $8\%$ for MRPC.

### v2

@@ 1 Introduction
In recent years, numerous PEFT approaches have been proposed [11, 12, 16, 17, 18, 30, 36], each varying in how and what they modify within FMs [10, 9]. These variations can be categorized by the position in the FMs where modifications are applied, the functions used for modification, and the methods of integrating the modifications. While effective, these methods require maintaining multiple adapter-like modules for each user-specific task on the FM server and the need to select and assemble a subset of these modules every time a batch of user requests is processed for inference [35] (see Fig. 1).
@@
  • We verify its effectiveness by evaluating it against full fine-tuning and state-of-the-art PEFT methods in nine tasks using seven different transformer backbones.
@@ 3 Proposed Methodology
In this section, we formally define the proposed approach, followed by an explanation of how it affects model learning as compared to existing soft-prompting approaches.
@@ 3.1 Preliminaries
Transformers. A transformer model consists of multiple stacked layers, where each layer has multiple heads ($=N_{H}$), each performing self-attention over the input [33]. Let us consider a single head, $H$ parameterised by the query, key, and value weights as $W^{Q},W^{K},W^{V}\in\mathcal{R}^{d_{H}\times d}$, respectively, where $d$ is the model dimension. In multi-headed attention, $d_{H}$ is typically set to $\frac{d}{N_{H}}$. For a given sequence of $n$ input vectors $\textbf{X}=\{\textbf{x}^{1},\ldots,\textbf{x}^{n}\}$ and a query vector $\textbf{x}^{i}$ with each $\textbf{x}\in\mathcal{R}^{d}$ , the output of the head at position $i$ is -
$\textbf{o}^{i}=\textrm{Attention}(W^{Q}\textbf{x}^{i},W^{K}\textbf{X},W^{V}\textbf{X})=\text{softmax}\big{(}W^{Q}\textbf{x}^{i}(W^{K}\textbf{X})^{\top}\big{)}W^{V}\textbf{X}$ (1)
where, we have ignored the constant $\sqrt{d_{H}}$ for notational convenience. For the first layer of the transformer network, the input X is the embedding matrix i.e. $\textbf{X}=\textbf{X}_{e}\in\mathcal{R}^{d\times n}$.
@@
Prompt-Tuning (PT). The objective of prompt-tuning is to achieve task-specific model adaptation by learning a "soft prompt" Z. Such a soft prompt is prepended to the input word embeddings and trained via back-propagation: $\max_{\textbf{Z}}\mathcal{L}=\mathbb{E}[\log p_{W}(\textbf{y}|\mathrm{concat}(\textbf{Z},\textbf{X}_{e}))]$. In PT, the soft prompt $\textbf{Z}\in\mathcal{R}^{d\times m}$ is parameterised by a set of $m$ learnable vectors $\{\textbf{z}^{1},\ldots\textbf{z}^{m}\}$, where $m$ denotes its length [15]. Thus, it can be interpreted as an embedding of virtual tokens that enable the model to perform a downstream task without updating its parameters [28, 8].
@@
where, $A_{ik}$ is the attention transformer gives to the prefix vector $\textbf{z}_{k}$ for a given query vector $\textbf{x}^{i}$. In Eq. 2, we can observe that the soft prompt linearly interpolates the head’s position-wise output; where the bias term can be considered in an offset subspace spanned by vectors $\{W^{V}\textbf{z}^{k}\}_{k=1}^{m}$ with dimension $m$ (or $\leq m$) [10].
3.2 Low-rank Prompt Adaptation (LoPA)
@@
Both $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$ have their own dedicated parameters. Similar to prompt tuning, $\textbf{Z}_{S}$ consists of $m$ learnable vectors. Conversely, $\textbf{Z}_{I}$ is obtained from input using the encoding function $f:\mathbb{R}^{d\times n}\to\mathbb{R}^{d\times m}$. However, encoding a matrix of size $d\times m$ can be expensive. For example, in [38], an MLP layer with hidden dimension $h$ requires $\mathcal{O}(hdm)$ parameters. To improve parameter efficiency, we assume a low-rank decomposition of $\textbf{Z}_{I}$ as $\textbf{Z}_{I}=\textbf{u}\times\textbf{v}^{\top}$ and encode the two matrices with rank $r$, $\textbf{u}\in\mathbb{R}^{d\times r}$ and $\textbf{v}\in\mathbb{R}^{m\times r}$, using separate MLP layers. This design makes $f$ computationally cheaper, with $\mathcal{O}(hdm(\frac{r}{d}+\frac{r}{m}))$ parameters, reduced by a factor of $(\frac{r}{d}+\frac{r}{m})<1$ when $r$ is chosen to be $\ll\min(d,m)$. The overall composition of Z can be represented as:
@@
where _ marks the three trainable components. The MLP head consists of a down- and up-projection layer with a non-linear activation in between. It sits atop a smaller language model, referred to as the Encoder model, that gives the input representation $\textbf{X}^{\prime}$. A visual illustration of the framework can be found in Fig. 2.
@@
Coupled learning of ${\textbf{Z}_{S}}$ and ${\textbf{Z}_{I}}$. Let’s examine the partial derivatives of the objective $\mathcal{L}$ with respect to $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$. Using the chain-rule on the formulation in Eq. 4 with $g(.)=\textrm{sigmoid}(.)$,
$\nabla_{\textbf{Z}_{S}}\mathcal{L}$ $=\nabla_{\textbf{Z}}\mathcal{L}\circ\sigma(\textbf{Z}_{I})$
$\nabla_{\textbf{Z}_{I}}\mathcal{L}$ $=(\nabla_{\textbf{Z}}\mathcal{L}\circ\textbf{Z}_{S})\cdot\sigma(\textbf{Z}_{I})(1-\sigma(\textbf{Z}_{I}))$
These expressions suggest a coupled learning process where changes in $\textbf{Z}_{I}$ (through the sigmoid function) directly impact the updates to $\textbf{Z}_{S}$, and the updates to $\textbf{Z}_{I}$ are scaled by both $\textbf{Z}_{S}$ and the derivative of the sigmoid function. In contrast, consider IDPG with $\textbf{Z}=\texttt{MLP}(\textbf{X}^{\prime})$ [38]. Here, the bias term of the MLP layer can be viewed as a task-specific element, such that the composition is $\textbf{Z}=\textbf{Z}_{S}+\textbf{Z}_{I}$. This approach results in updates to $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$ being independent of each other. Such a linear-sum operation may fail to capture the complex relationships between the two components that LoPA can capture due to LoPA’s non-linear factorization.
@@ 4.1 Experimental Setup
Tasks. We evaluate LoPA on (i) six Natural Language Understanding (NLU) tasks from the GLUE benchmark [34]—namely, SST-2 [31], MNLI [37], MRPC [3], QNLI [29], QQP, and RTE [5]; (ii) a code-generation task that requires the model to complete method bodies from MBPP benchmark [1], and (iii) two code-understanding tasks—namely, CruxEval-I (input prediction) and CruxEval-O (output prediction) from CruxEval benchmark [6]. These tasks assess the model’s ability to reason about the execution of simple Python programs by asking it to predict the input given the output of a function and vice-versa.
Baselines. We evaluate LoPA against the following baselines that represent different customisation approaches. (1) FFT, (2) LoRA [12], (3) p-tuning-v2 [20] and (4) prefix-tuning (h=512) [16] are selected as representatives of methods that customise models on the server side, with FFT representing full fine-tuning and the remainder showcasing parameter-efficient techniques. (5) Standard prompt-tuning [15], (6) S-IDPG [38] and (7) DePT [30] are chosen because they customise models from the user side, focusing on soft prompt insertion with DePT also learning updates to the input embedding matrix. For IDPG, we use DNN layers in MLP-head to encode the soft prompt for the input embedding layer. We chose DNN over PHM layers [39] because we found them to exhibit better performance across our task set. Refer to Appendix 7.1 for the comparison. Additionally, the proposed LoPA also uses DNN layers, facilitating a fair comparison with S-IDPG-DNN [38]. Lastly, we include results for the model without any fine-tuning to report its (8) zero-shot performance.
Backbone Architectures. For NLU tasks, we test all the tuning methods on 355M RoBERTa [22] similar to the setup opted by prior work [20, 38, 41]. For Code Generation and Understanding, we test a subset of the baselines (FFT, LoRA, PT, S-IDPG) on a range of FM backbones; starting from smaller FMs: 350M CodeGen-mono [25] and 1.3B Deepseek-Coder [7]; mid-sized FMs: 2.7B phi-2 [13], 3.8B phi-3 [24]; and larger FMs: 7B Deepseek-Coder [7] and 8B Llama 3 [23].
@@
Training Configuration. For NLU tasks, training with FFT and LoRA was done for 10 epochs, while with prompt-tuning-based approaches it was done for 20 epochs. In MBPP, all foundation model (FM) backbones were trained for 10 epochs across all tuning methods. In CruxEval Tasks across all PEFT methods, FM backbones under 7B were trained for 20 epochs, while larger FMs ($\geq$7B) were trained for 10 epochs. Lastly, training with FFT on CruxEval tasks was done for 5 epochs. The learning rates for LoPA are set to $1\times 10^{-5}$ in NLU and $1\times 10^{-3}$ in Coding tasks. The baseline tuning methods use the following learning rates across all the tasks: FFT using $1\times 10^{-5}$, LoRA and the remainder of soft-prompting approaches using $1\times 10^{-4}$. All experiments are conducted on 40GB 2xA100 GPUs.
Evaluation. We report binary or multi-class classification accuracies and F1 scores for NLU tasks as provided in the GLUE benchmark. For coding tasks, we report the $\textit{pass}@1$ scores computed using the best-performing temperatures: $0.6$ for MBPP and $0.2$ for CruxEval.
@@ 4.2 Baseline Comparison
Performance on Natural Language Understanding. In Table 1, we can observe that the LoPA consistently outperforms the traditional prompt-tuning method and DePT by an average margin of $28.62$ points and $25.39$ points respectively. This result demonstrates that conditioning the soft prompt on instances enables it to influence the model’s output more significantly. Furthermore, LoPA shows an average improvement of $5.36$ points over IDPG, highlighting that the proposed factorisation captures complex relationships between task-specific ($\textbf{Z}_{S}$) and instance-specific ($\textbf{Z}_{I}$) components. This improvement is particularly notable in limited-data settings, with a $12.5$-point increase in MRPC and a $6.13$-point increase in RTE.
@@
[…]
(acc) (acc) (acc, F1) (acc) (acc, F1) (acc)
$\textrm{concat}(\textbf{Z}_{S},\textbf{Z}_{I})$ 94.50 78.45 76.00 91.43 76.11 84.12 83.44
$\max(\textbf{Z}_{S},\textbf{Z}_{I})$ 95.99 89.37 88.62 93.74 78.44 85.92 88.68
$\textbf{Z}_{S}\circ g(\textbf{Z}_{I})$ 95.99 89.22 91.09 93.74 89.72 83.39 90.53
[…]
Performance on Code Understanding. In Table 2, LoPA consistently improves the $\textit{pass}@1$ score of the baseline with no tuning across all FM backbones. It outperforms prompt-tuning on CruxEval tasks, with modest improvements of approximately $2$ to $4$ points on smaller FMs like CodeGen and DeepSeekCoder-1.3B and larger improvements ranging from $8$ to $11$ points on larger FMs like LLama-3 and Phi-3 in CruxEval-O. Furthermore, IDPG performs worse than PT for all models except Phi-3 in CruxEval-O. These results suggest that merely encoding an instance-sensitive soft prompt does not guarantee improvements and can even degrade performance (e.g., IDPG on CodeGen and DeepSeekCoder-1.3B in CruxEval tasks). The poor generalization of the learned soft prompt, possibly due to over-parameterization and resulting overfitting, might explain this behaviour. In contrast, LoPA, being more parameter-efficient, explicitly incorporates task and instance information in its design of the soft prompt, leading to better performance.
LoPA also performs on par with LoRA, often within a range of $1$ to $4$ points of $\textit{pass}@1$, while roughly using two-thirds of the parameters. Notably, LoPA outperforms LoRA across all models in CruxEval-I, except for LLama-3, with improvements approximately ranging from $2$ points in DeepSeekCoder-7B to $4$ points in Phi-3. This result might be attributed to the fact that many of the FMs considered here are good knowledge approximators, well-trained on diverse datasets, and demonstrate strong zero-shot generalization. Directly updating a subset of their weights can still lead to catastrophic forgetting, where the models lose previously acquired knowledge [27]. In such cases, soft prompting, as employed by LoPA, can effectively elicit the necessary skills to solve new tasks without compromising existing knowledge [26].
Performance on Code Completion. On the code completion task of MBPP, both IDPG and LoPA improve the performance of the baseline model with nearly equal gains except for LLama-3. However, LoPA achieves this with significantly fewer tunable parameters—approximately half the number used in CodeGen and only one-fifth of those used by IDPG in DeepseekCoder-7b. This demonstrates that LoPA scales well with the size of the foundation model, maintaining both performance and parameter efficiency. This efficiency is attributed to the low-rank approximation of the instance-specific matrix employed by LoPA. For LLama-3, all tuning approaches led to a drop in performance, possibly due to over-fitting, suggesting that LLama-3 might already be trained on MBPP.
@@ 4.3 Ablation Study
Performance of LoPA as a function of soft-prompt length. In Figure 3, we study how the length of the soft prompt impacts the performance of LoPA compared to other prompting methods. Increasing the length of the soft prompt corresponds to adding more vectors to the set that represents the soft prompt, thus expanding the offset subspace (See Eq. 2). Whether the added vectors are mutually independent and provide additional useful information depends on the tuning approach to learning them and the offset subspace of the FM.
@@
Performance of LoPA as a function of rank. In Figure 4, we examine how the rank of the proposed low-rank decomposition of the instance-specific component affects the performance. For NLU, we consider three tasks: MRPC, RTE, and SST-2, and observe the performance of RoBERTa as the rank increases from 1 to 4 . Performance consistently improves with increasing rank, showing gains of $1\%$ to $2\%$ for SST-2 and up to $8\%$ for MRPC.
@@
Performance of LoPA with respect to function encoding Z. In Table 3, we evaluate the effect of different functions encoding Z as a combination of $\textbf{Z}_{I}$ and $\textbf{Z}_{S}$ on RoBERTa’s performance in NLU tasks. The results show that simply concatenating $\textbf{Z}_{S}$ and $\textbf{Z}_{I}$ performs the worst while the non-linear functions, such as max(.) and the proposed gating mechanism, yield the best performance. We leave the exploration of other non-linear functions for future work.